OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6? Answer x + y = 7 x + y = 1 x – y = –7 x – y = –1

OpenStudy (cwrw238):

slope = -1 use standard form y-y1 = m(x-x1) where m = -1, x1 = 3 and y1 = 4

OpenStudy (hoblos):

hope you got it by yourself anyway the answer is x + y = 7

OpenStudy (anonymous):

I JUST DID THE FORMULA THANKSS